# Computation of Robust Option Prices via Structured Multi-Marginal Martingale Optimal Transport

### v1

@@ 1 Introduction
In this article, we combine entropic regularisation with the exploitation of structures inherent in the problem, in order to computationally solve the multi-marginal MOT problem (2). In line with the above-mentioned literature on graph-structured multi-marginal OT problems, the idea is to exploit specific structures in the cost function. We therefore restrict ourselves to financial derivatives whose payoff is of the form
@@
where $X_{0}=h_{0}(S_{0})$ and $X_{t}=h_{t}(S_{t},S_{t-1},X_{t-1})$, for some functions $h_{0}:\mathbb{R}\to\mathbb{R}$, $h_{t}:\mathbb{R}^{3}\to\mathbb{R}$ and $\phi_{t}:\mathbb{R}^{4}\to\mathbb{R}$, $t\in\{1,\dots,T\}$. We utilise this structure in a twofold manner. First, leveraging on the fact that for a given stochastic process, one can construct a Markov process with the same pair-wise joint marginal distributions onto adjacent marginals, for payoffs of the form (3), we manage to relax condition (2b) to $\mathbb{E}_{\mathbb{Q}}[S_{t}|\sigma(S_{t-1},X_{t-1})]=S_{t-1}$, $t\in\{1,\dots,T\}$. In effect, we are facing a problem where each constraint is only connecting adjacent marginals and where the cost function has the same structure as in the above-mentioned works on graph-structured problems. In turn, we may thus employ methods from the theory on graph-structured multi-marginal OT problems to find an efficient algorithm for solving problem (2). Our first main result establishes the equivalence between the relaxed and the original problem formulation, while our second main result provides an efficient method for solving the corresponding entropy regularised problem. We illustrate the usefulness of our method by performing explicit computations in a number of examples: First, we apply our algorithm to problems for which the theoretical solution is known; we use it to recover the left-monotone coupling, late and early transport behaviour, as well as optimal models for digital options and lookback options. We then use it to compute the robust price bound for an Asian option, subject to several marginal constraints, where as far as we know, no theoretical solution is known.
For MOT problems, entropic regularisation was first introduced by de March [18], who derived a method based on Sinkhorn’s algorithm and Newton’s method for solving bi-marginal problems. As for early numerical results on robust pricing problems, we also mention Henry-Labordere [34], Davis, Obłój and Raval [17] and Tan and Touzi [59]. Computationally solving multi-marginal MOT problems has since been addressed with neural network approaches in Eckstein and Kupper [21] and Eckstein et al. [20], and by linear programming approaches in [20] and Guo and Obłój [29], where the latter relies on a relaxation of the martingale constraint. However, due to the curse of dimensionality, it is computationally challenging to utilise these methods for problems with a large number of marginals. Markovian assumptions in a multi-marginal MOT context were introduced in Henry-Labordère [37, Remark 2.7] and further discussed in Sester [54], who considered a problem subject to an explicit Markov constraint. Such problems are not convex, which motivated Eckstein and Kupper [22] to consider multi-marginal MOT problems under a time-homogeneity assumption. While Alfonsi et al. [1] and Lindheim and Steidl [61] have developed methods for OT problems featuring constraints closely related to the martingale constraint, to the best of our knowledge, there are no other works using entropic regularisation methods for multi-marginal MOT problems specifically.
@@
$\langle A,B\rangle=\sum_{i_{0},\dots,i_{n}\in[m]}A(i_{0},\dots,i_{n})B(i_{0},\dots,i_{n}),\quad A,B\in\mathbb{R}^{n^{m}},$
@@
$P_{j}(A)(i_{j}):=\sum_{i_{\ell}:\ell\in[m]\backslash j}A(i_{0},\dots i_{m}),\quad i_{j}=1,\dots,n,\quad\mbox{ where }A\in\mathbb{R}^{n^{m}}.$
Similarly, we let $P_{j,k}:\mathbb{R}^{n^{m}}\rightarrow\mathbb{R}^{n^{2}}$ be the projection operator that projects the tensor jointly onto the two dimensions $j$ and $k$ (see [23, p. 7]). Every real-valued function is assumed to be Borel measurable.
@@ 2.1 Problem formulation for structured payoff functions
where $h_{0}:\mathbb{R}\rightarrow\mathbb{R}$ and $h_{t}:\mathbb{R}^{3}\rightarrow\mathbb{R}$ for $t\in[T]\backslash 0$. By defining $X$ in this way we are able to isolate important features of the path of $S$, thus enabling the pricing of many exotic options. We are here interested in pricing path-dependent derivatives whose payoff function can be pairwise decoupled, i.e.,
@@
Here we assume that the given marginals $\mu_{t}$, for $t\in\mathcal{T}\subseteq\{0,1,\dots,T\}$, are in convex order. Such problems are well posed, since it follows by Strassen’s theorem that their feasible sets are nonempty [58].
@@ 2.2 Financial examples
the maximum process is of the form (4) (the minimum process can be considered analogously). Related derivatives are single asset, single period maximum lookback options such as the floating strike lookback put, whose payoff function is $\phi(S_{T},X_{T})=\left(X_{T}-S_{T}\right)^{+}$ (see, for example, [39, pp. 623–525]). Certain barrier options can also be priced by considering the maximum process, for example the up-and-in barrier call with barrier $b$ and strike $K$, which at maturity pays out $\phi(S_{T},X_{T})=\chi_{[b,\infty)}(X_{T})(S_{T}-K)^{+}$ (see, for example, [12, pp. 143–150]). In Section 5 we will solve an upper bound MOT problem where the payoff is equal to the terminal value of the maximum.
@@
it is of the form (4). Derivatives whose payoff is a function of the arithmetic mean are Asian options with arithmetic averaging, such as the average price call and the average strike call. Their payoffs are $\phi(X_{T})=(X_{T}-K)^{+}$ and $\phi(S_{T},X_{T})=(S_{T}-X_{T})^{+}$, respectively, where $K>0$ denotes a fixed strike price (see, for example, [39, pp. 626–627]). In Section 5 we will provide an example where we compute the robust fair price of an Asian straddle, given two, three and four known marginals.
@@
Its terminal value, $X_{T}$, indicates whether the price $S$ has passed through all the sets $A_{0},\dots,A_{T}$ during the duration of the contract. By setting all but two of the sets $A_{0},\dots,A_{T}$ equal to $\mathbb{R}$, it is possible to consider second-order binary options. Indeed, let $T_{0}\in[T-1]\backslash 0$ and define $A_{T_{0}}:=\{x\in\mathbb{R}:x\geq b_{0}\}$ and $A_{T}:=\{x\in\mathbb{R}:x\geq b\}$ for some barrier levels $b_{0}$ and $b$. Let $A_{t}=\mathbb{R}$ for $t\in[T]\backslash\{T_{0},T\}$. Then the payoff of an up-up type second-order asset binary option can be written $\phi(S_{T},X_{T})=S_{T}X_{T}$ (see, for example, [12, pp. 109–110]). A double knockout option, on the other hand, expires worthless if the price of the underlying crosses any of the two barriers $b_{l}<S_{0}<b_{u}$ (see, for example, [49, p. 274]). Defining $A:=\{x\in\mathbb{R}:b_{l}<x<b_{u}\}$ and setting $A_{t}=A$ for all $t\in[T]$ allows us to also consider such options. In Section 5 we will compute the robust price of a digital option.
@@
Consider the case when the process $X$ is the realised variance of the price process $S$, that is, when $X_{t}=t^{-1}\sum_{j=1}^{t}(\log(S_{j}/S_{j-1}))^{2}$ for $t\in[T]\backslash[0]$. Set $X_{0}:=0$; then $X$ can be written as
$X_{t}=\begin{cases}0,&t=0\\ \left(1-\frac{1}{t}\right)X_{t}+\frac{1}{t}\left(\log\left(\frac{S_{t}}{S_{t-1}}\right)\right)^{2},&t\in[T]\backslash 0,\end{cases}$
which shows that the realised variance is indeed of the form (4). The payoff of a variance swap is then given by $\phi(X_{T})=X_{T}-\sigma_{\text{fixed}}^{2}$, where $\sigma_{\text{fixed}}^{2}$ denotes the fixed variance (see, for example, [39, p. 629]). Needless to say, the realised volatility of the price $S$ can easily be obtained as $\sqrt{X_{t}}$ at any time point $t\in[T]$. We conclude this example by noting that the optimal solution of the corresponding MOT problem is known in the case when $T=1$; it is the left-monotone transport plan [7, 36]. We will revisit this solution in Section 5.
@@
and its terminal value, $X_{T}$, corresponds to the number of, say, days that the price $S$ has spent within $A$ during the duration of the contract, a property that allows for pricing Parisian options. In order for such an option — written with a $D\in\mathbb{N}$ day window — to activate, the price of the underlying must spend at least $D$ days within $A$ (see, for example, [39, p. 622]). As an example, an up-and-in Parisian put with a $D$ day window, a strike price $K$ and a barrier $b$ has a payoff that can be written $\phi(S_{T},X_{T})=(K-S_{T})^{+}\chi_{[D,\infty)}(X_{T})$, with $X_{T}$ defined as above for $A=\{x\in\mathbb{R}:x\geq b\}$.
@@
which is of the form (4). This choice of $X$ allows for pricing dual expiry options — the payoff of a forward start call is given by $\phi(S_{T},X_{T})=(S_{T}-X_{T})^{+}$, while a ratchet call option with strike price $K$ has a payoff $\phi(S_{T},X_{T})=\max\left\{(S_{T}-K)^{+},(X_{T}-K)^{+}\right\}$ (see, for example, [12, pp. 107, 119]).
@@
The payoff function of a globally floored, locally capped cliquet option is then given by $\phi(X_{T})=\max\{X_{T},F_{g}\}$, where $F_{g}$ is the global floor (see, for example, [62, p. 379]).
@@ 2.3 Structured multi-marginal optimal transport
The OT problem was introduced in Section 1 as the problem of minimising the total cost for transporting the mass of one distribution to another one. The transport can occur in one or several steps — we refer to the former as bi-marginal transport, while the latter is multi-marginal transport. In order to formalise this, consider $m+1$ probability spaces $(\mathcal{X}_{0},\mathcal{F}_{\mathcal{X}_{0}},\mu_{0}),\dots,(\mathcal{X}_{m},\mathcal{F}_{\mathcal{X}_{m}},\mu_{m})$ and let $\mathcal{M}(\mathcal{X}_{j})$ denote the set of probability measures on $\mathcal{X}_{j}$ for $j\in[m]$. Similarly, let $\mathcal{M}(\mathcal{X})$ denote the set of probability measures on the product space $\mathcal{X}:=\mathcal{X}_{0}\times\dots\times\mathcal{X}_{m}$ equipped with the product $\sigma$-algebra. Let the projection operators $\mathcal{P}_{j}:\mathcal{M}(\mathcal{X})\rightarrow\mathcal{M}(\mathcal{X}_{j})$ be defined as $\mathcal{P}_{j}(\pi):=\int_{\mathcal{X}_{0}}\dots\int_{\mathcal{X}_{j-1}}\int_{\mathcal{X}_{j+1}}\dots\int_{\mathcal{X}_{m}}\mathrm{d}\pi$, for $\pi\in\mathcal{M}(\mathcal{X})$ and $j\in[m]$. A measure $\pi\in\mathcal{M}(\mathcal{X})$ is said to be a coupling of $(\mu_{0},\dots,\mu_{m})$ if it satisfies $\mathcal{P}_{j}(\pi)=\mu_{j}$ for $j\in[m]$; in the context of optimal transport, a coupling is often referred to as a transport plan. Given a cost function $c:\mathcal{X}\rightarrow\mathbb{R}$, where $c(x_{0},\dots,x_{m})$ is the cost associated with the path $(x_{0},\dots,x_{m})\in\mathcal{X}$, the optimal transport problem is the problem of finding a coupling that minimises the total associated cost, i.e. that attains
$\underset{\pi\in\mathcal{M}(\mathcal{X})}{\inf}\quad$ $\int_{\mathcal{X}_{0}\times\dots\times\mathcal{X}_{m}}c(x_{0},\dots,x_{m})\mathrm{d}\pi(x_{0},\dots,x_{m})$ (7a)
subject to $\mathcal{P}_{j}(\pi)=\mu_{j},\quad j\in[m].$ (7b)
@@
A coupling of $(\mu_{0},\dots,\mu_{m})$ that attains the minimum in problem (7) is said to be an optimal coupling or an optimal transport plan. Problem (7) can be generalised by relaxing the constraint (7b) to only hold for a subset $\Gamma\subseteq[m]$ of the marginals or by introducing unbalanced problems [4, 14, 23] (cf. bi-marginal problems [28, 42, 43, 52]).
@@
In order to solve problems on the form (7) computationally, we assume that each of the sets $\mathcal{X}_{0},\dots,\mathcal{X}_{m}$ is finite, that is, $\cup_{j\in[m]}\mathcal{X}_{j}$ is contained within $n\in\mathbb{N}$ points. Then the marginals can be represented by vectors $\upmu_{j}\in\mathbb{R}^{n}$, $j\in[m]$, and the cost function and the transport plan by tensors $C\in\mathbb{R}^{n^{m+1}}$ and $Q\in\mathbb{R}_{+}^{n^{m+1}}$, respectively. The discrete analogue to problem (7) is then
$\underset{Q\in\mathbb{R}_{+}^{n^{m+1}}}{\min}\quad$ $\langle C,Q\rangle$ (8a)
subject to $P_{j}(Q)=\upmu_{j},\quad j\in[m].$ (8b)
@@
Problem (8) is a linear programming problem in $n^{m+1}$ variables. Since the number of variables grows exponentially as a function of the number of marginals, the problem quickly becomes intractable for standard methods — even problems with two marginals ($m=2$) are not straightforward to solve when the number of points $n$ is large. One way to alleviate this difficulty is to regularise the problem with an entropy term, as proposed in [8, 16, 51]; we will see that this leads to a dual problem in $n(m+1)$ variables and explicit formulas for the optimal choice thereof.
@@
$\underset{Q\in\mathbb{R}_{+}^{n^{m+1}}}{\min}\quad$ $\langle C,Q\rangle+\varepsilon D(Q)$ (9a)
subject to $P_{j}(Q)=\upmu_{j},\quad j\in[m],$ (9b)
@@
$D(Q):=\sum_{i_{0},\dots,i_{m}\in[n]}\left(Q(i_{0},\dots,i_{m})\log Q(i_{0},\dots,i_{m})-Q(i_{0},\dots,i_{m})+1\right)$
@@
$\mathcal{L}(Q,\lambda)=\langle C,Q\rangle+\varepsilon D(Q)+\sum_{j\in[m]}\lambda_{j}^{\top}\left(\upmu_{j}-P_{j}(Q)\right),$
@@
where $\lambda=\{\lambda_{j}\}_{j\in[m]}$ are Lagrangian multipliers in $\mathbb{R}^{n}$ corresponding to the constraints (9b). The Lagrangian multipliers are also referred to as dual variables, while the dual functional of problem (9) is defined as $\varphi(\lambda):=\min\{\mathcal{L}(Q,\lambda):Q\in\mathbb{R}_{+}^{n^{m+1}}\}$.
@@
Since the regularisation term $D(Q)$ ensures that the Lagrangian function is strictly convex in $Q$, the minimum is attained by the unique $Q>0$ for which the gradient of the Lagrangian function is zero; that is, for the transport plan $Q$ satisfying
$0=\partial_{Q(i_{0},\ldots,i_{m})}\mathcal{L}(Q,\lambda)=C(i_{0},\ldots,i_{m})+\varepsilon\log Q(i_{0},\ldots,i_{m})-\sum_{j\in[m]}\lambda_{j}(i_{j})$
@@
for $i_{0},\dots,i_{m}\in\{1,\dots,n\}$. Solving for $Q$ yields that the primal solution is given by $Q_{\lambda}=K\odot U_{\lambda}$, where
$K(i_{0},\dots,i_{m}):=e^{-C(i_{0},\dots,i_{m})/\varepsilon},\quad U_{\lambda}(i_{0},\dots,i_{m}):=\prod_{j\in[m]}e^{\lambda_{j}(i_{j})/\varepsilon},\quad i_{0},\dots,i_{m}=1,\dots,n$

### v2

@@ 1 Introduction
For MOT problems, entropic regularisation was first introduced by de March [21], who derived a method based on Sinkhorn’s algorithm and Newton’s method for solving bi-marginal problems. As for early numerical results on robust pricing problems, we also mention Henry-Labordere [38], Davis, Obłój and Raval [20] and Tan and Touzi [66]. Computationally solving multi-marginal MOT problems has since been addressed with neural network approaches in Eckstein and Kupper [24] and Eckstein et al. [23], and by linear programming approaches in [23] and Guo and Obłój [32], where the latter relies on a relaxation of the martingale constraint. However, due to the curse of dimensionality, it is computationally challenging to utilise these methods for problems with a large number of marginals. Markovian assumptions in a multi-marginal MOT context were introduced in Henry-Labordère [41, Remark 2.7] and further discussed in Sester [60], who considered a problem subject to an explicit Markov constraint. Such problems are not convex, which motivated Eckstein and Kupper [25] to consider multi-marginal MOT problems under a time-homogeneity assumption. Other related works include Alfonsi et al. [1] and Lindheim and Steidl [68], who have developed methods for OT problems featuring moment and affine constraints, respectively.
To the best of our knowledge, there are no previous methods that allows for computing multi-marginal MOT problems with a large number of marginals. This article aims to fill this gap, and a main contribution is a method that can efficiently compute solutions to such problems. We combine entropic regularisation with the exploitation of structures inherent in the problem, in order to computationally solve the multi-marginal MOT problem (2). In line with the above-mentioned literature on graph-structured multi-marginal OT problems, the idea is to exploit specific structures in the cost function. However, the presence of the martingale constraint introduces dependencies between the marginals and as a result, a multi-marginal MOT problem is not sparse — this is a fundamental difficulty when it comes to addressing such problems computationally. To get around this problem we first restrict ourselves to financial derivatives whose payoff is of the form
@@
where $X_{0}=h_{0}(S_{0})$ and $X_{t}=h_{t}(S_{t},S_{t-1},X_{t-1})$, for some functions $h_{0}:\mathbb{R}\to\mathbb{R}$, $h_{t}:\mathbb{R}^{3}\to\mathbb{R}$ and $\phi_{t}:\mathbb{R}^{4}\to\mathbb{R}$, $t\in\{1,\dots,T\}$. We then utilise this structure in a twofold manner. First, leveraging on the fact that for a given stochastic process, one can construct a Markov process with the same pair-wise joint marginal distributions onto adjacent marginals, for payoffs of the form (3), we manage to relax condition (2b) to $\mathbb{E}_{\mathbb{Q}}[S_{t}|\sigma(S_{t-1},X_{t-1})]=S_{t-1}$, $t\in\{1,\dots,T\}$. Note that this is done without introducing any explicit Markovian assumption on the price process or any assumption regarding the convexity of the payoff function. In effect, we are facing a problem where each constraint is only connecting adjacent marginals and we may thus employ methods from the theory on graph-structured multi-marginal OT problems to find an efficient algorithm for solving problem (2). We mention that we, in one of our numerical examples (see Section 5.2), address an MOT problem with more than 50 marginals, where each marginal is supported on a large number of points. To the best of our knowledge, this is significantly larger than what has been reported for other methods. We illustrate the usefulness of our method by applying it on a number of examples: First, we apply our algorithm to problems for which the theoretical solution is known; we use it to recover the left-monotone coupling, late and early transport behavior, as well as optimal models for digital options and lookback options. We then use it to compute the robust price bound for an Asian option, subject to several marginal constraints, where as far as we know, no theoretical solution is known.
@@
$\langle\mathbf{A},\mathbf{B}\rangle=\sum_{i_{0},\dots,i_{n}\in[m]}\mathbf{A}(i_{0},\dots,i_{n})\mathbf{B}(i_{0},\dots,i_{n}),\quad\mathbf{A},\mathbf{B}\in\mathbb{R}^{n^{m}},$
@@
$P_{j}(\mathbf{A})(i_{j}):=\sum_{i_{\ell}:\ell\in[m]\backslash j}\mathbf{A}(i_{0},\dots i_{m}),\quad i_{j}=1,\dots,n,\quad\mbox{ where }\mathbf{A}\in\mathbb{R}^{n^{m}}.$
Similarly, we let $P_{j,k}:\mathbb{R}^{n^{m}}\rightarrow\mathbb{R}^{n^{2}}$ be the projection operator that projects the tensor jointly onto the two dimensions $j$ and $k$ (see [26, p. 7]).
Every real-valued function is assumed to be Borel measurable and we let the regular conditional distribution be equal to some real-valued constant if conditioning on a null event. Zero is assumed to be included in the set of positive real numbers and we use the convention of $0\cdot\infty=0$.
@@ 2.1 Problem formulation for structured payoff functions
where $h_{0}:\mathbb{R}\rightarrow\mathbb{R}$ and $h_{t}:\mathbb{R}^{3}\rightarrow\mathbb{R}$ for $t\in[T]\backslash 0$. By defining $X$ in this way we are able to isolate important features of the path of $S$, thus enabling the pricing of many exotic options. We are here interested in pricing path-dependent derivatives whose payoff function $\phi:\mathbb{R}^{T+1}\rightarrow\mathbb{R}$ is finite and can be pairwise decoupled, that is,
@@
Here we assume that the given marginals $\mu_{t}$, for $t\in\mathcal{T}\subseteq\{0,1,\dots,T\}$, are in convex order. Such problems are well posed, since it follows by Strassen’s theorem that their feasible sets are nonempty [64]. We remark that problems on the above given form are convex, albeit not strictly convex, in $\mathbb{Q}$; however, showing the existence of optimal solutions requires further assumptions on the payoff function which we do not want to introduce here (see, for example, [6, Theorem 1.1]).
@@
Remark 2.1.
We here limit ourselves to the case when the auxiliary process $X$ takes values on $\mathbb{R}$, but it is our belief that our framework could be extended to the case when the process $X$ takes values on $\mathbb{R}^{d}$ for $d>1$. Note that this would allow us to take more information about the path of the price process $S$ into account, but only at the cost of the size of the problem increasing — we emphasise that our computational framework is useful only when $d$ is significantly smaller than $T$.
@@ 2.2 Financial examples
the maximum process is of the form (4) (the minimum process can be considered analogously). Related derivatives are single asset, single period maximum lookback options such as the floating strike lookback put, whose payoff function is $\phi(S_{T},X_{T})=\left(X_{T}-S_{T}\right)^{+}$ (see, for example, [45, pp. 623–525]). Certain barrier options can also be priced by considering the maximum process, for example the up-and-in barrier call with barrier $b$ and strike $K$, which at maturity pays out $\phi(S_{T},X_{T})=\chi_{[b,\infty)}(X_{T})(S_{T}-K)^{+}$ (see, for example, [14, pp. 143–150]). In Section 5 we will solve an upper bound MOT problem where the payoff is equal to the terminal value of the maximum.
@@
it is of the form (4). Derivatives whose payoff is a function of the arithmetic mean are Asian options with arithmetic averaging, such as the average price call and the average strike call. Their payoffs are $\phi(X_{T})=(X_{T}-K)^{+}$ and $\phi(S_{T},X_{T})=(S_{T}-X_{T})^{+}$, respectively, where $K>0$ denotes a fixed strike price (see, for example, [45, pp. 626–627]). In Section 5 we will provide an example where we compute the robust fair price of an Asian straddle, given two, three and four known marginals.
@@
Its terminal value, $X_{T}$, indicates whether the price $S$ has passed through all the sets $A_{0},\dots,A_{T}$ during the duration of the contract. By setting all but two of the sets $A_{0},\dots,A_{T}$ equal to $\mathbb{R}$, it is possible to consider second-order binary options. Indeed, let $T_{0}\in[T-1]\backslash 0$ and define $A_{T_{0}}:=\{x\in\mathbb{R}:x\geq b_{0}\}$ and $A_{T}:=\{x\in\mathbb{R}:x\geq b\}$ for some barrier levels $b_{0}$ and $b$. Let $A_{t}=\mathbb{R}$ for $t\in[T]\backslash\{T_{0},T\}$. Then the payoff of an up-up type second-order asset binary option can be written $\phi(S_{T},X_{T})=S_{T}X_{T}$ (see, for example, [14, pp. 109–110]). A double knockout option, on the other hand, expires worthless if the price of the underlying crosses any of the two barriers $b_{l}<S_{0}<b_{u}$ (see, for example, [55, p. 274]). Defining $A:=\{x\in\mathbb{R}:b_{l}<x<b_{u}\}$ and setting $A_{t}=A$ for all $t\in[T]$ allows us to also consider such options. In Section 5 we will compute the robust price of a digital option.
@@
Consider the case when the process $X$ is the realised variance of the price process $S$, that is, when $X_{t}=t^{-1}\sum_{j=1}^{t}(\log(S_{j}/S_{j-1}))^{2}$ for $t\in[T]\backslash 0$. Set $X_{0}:=0$; then $X$ can be written as
$X_{t}=\begin{cases}0,&t=0\\ \big{(}1-\frac{1}{t}\big{)}X_{t-1}+\frac{1}{t}\big{(}\log\big{(}\frac{S_{t}}{S_{t-1}}\big{)}\big{)}^{2},&t\in[T]\backslash 0,\end{cases}$
which shows that the realised variance is indeed of the form (4). The payoff of a variance swap is then given by $\phi(X_{T})=X_{T}-\sigma_{\text{fixed}}^{2}$, where $\sigma_{\text{fixed}}^{2}$ denotes the fixed variance (see, for example, [45, p. 629]). Needless to say, the realised volatility of the price $S$ can easily be obtained as $\sqrt{X_{t}}$ at any time point $t\in[T]$. We note that the optimal solution of the corresponding MOT problem is known in the case when $T=1$; it is the left-monotone transport plan [7, 40], a solution that will be revisited in Section 5.
@@
and its terminal value, $X_{T}$, corresponds to the number of, say, days that the price $S$ has spent within $A$ during the duration of the contract, a property that allows for pricing Parisian options. In order for such an option — written with a $D\in\mathbb{N}$ day window — to activate, the price of the underlying must spend at least $D$ days within $A$ (see, for example, [45, p. 622]). As an example, an up-and-in Parisian put with a $D$ day window, a strike price $K$ and a barrier $b$ has a payoff that can be written $\phi(S_{T},X_{T})=(K-S_{T})^{+}\chi_{[D,\infty)}(X_{T})$, with $X_{T}$ defined as above for $A=\{x\in\mathbb{R}:x\geq b\}$.
@@
which is of the form (4). This choice of $X$ allows for pricing dual expiry options — the payoff of a forward start call is given by $\phi(S_{T},X_{T})=(S_{T}-X_{T})^{+}$, while a ratchet call option with strike price $K$ has a payoff $\phi(S_{T},X_{T})=\max\left\{(S_{T}-K)^{+},(X_{T}-K)^{+}\right\}$ (see, for example, [14, pp. 107, 119]).
@@
The payoff function of a globally floored, locally capped cliquet option is then given by $\phi(X_{T})=\max\{X_{T},F_{g}\}$, where $F_{g}$ is the global floor (see, for example, [69, p. 379]).
@@ 2.3 Structured multi-marginal optimal transport
The OT problem was introduced in Section 1 as the problem of minimising the total cost for transporting the mass of one distribution to another one. The transport can occur in one or several steps — we refer to the former as bi-marginal transport, while the latter is multi-marginal transport. In order to formalise this, consider $T+1$ probability spaces $(\mathcal{X}_{0},\mathcal{F}_{\mathcal{X}_{0}},\mu_{0}),\dots,(\mathcal{X}_{T},\mathcal{F}_{\mathcal{X}_{T}},\mu_{T})$ and let $\mathcal{M}(\mathcal{X}_{j})$ denote the set of probability measures on $\mathcal{X}_{t}$ for $t\in[T]$. Similarly, let $\mathcal{M}(\mathcal{X})$ denote the set of probability measures on the product space $\mathcal{X}:=\mathcal{X}_{0}\times\dots\times\mathcal{X}_{T}$ equipped with the product $\sigma$-algebra. Let the projection operators $\mathcal{P}_{t}:\mathcal{M}(\mathcal{X})\rightarrow\mathcal{M}(\mathcal{X}_{t})$ be defined as $\mathcal{P}_{t}(\pi):=\int_{\mathcal{X}_{0}}\dots\int_{\mathcal{X}_{t-1}}\int_{\mathcal{X}_{t+1}}\dots\int_{\mathcal{X}_{T}}\mathrm{d}\pi$, for $\pi\in\mathcal{M}(\mathcal{X})$ and $t\in[T]$. A measure $\pi\in\mathcal{M}(\mathcal{X})$ is said to be a coupling of $(\mu_{0},\dots,\mu_{T})$ if it satisfies $\mathcal{P}_{t}(\pi)=\mu_{t}$ for $t\in[T]$; in the context of optimal transport, a coupling is often referred to as a transport plan. Given a cost function $c:\mathcal{X}\rightarrow\mathbb{R}$, where $c(x_{0},\dots,x_{T})$ is the cost associated with the path $(x_{0},\dots,x_{T})\in\mathcal{X}$, the optimal transport problem is the problem of finding a coupling that minimises the total associated cost, i.e. that attains
$\underset{\pi\in\mathcal{M}(\mathcal{X})}{\inf}\quad$ $\int_{\mathcal{X}_{0}\times\dots\times\mathcal{X}_{T}}c(x_{0},\dots,x_{T})\mathrm{d}\pi(x_{0},\dots,x_{T})$ (7a)
subject to $\mathcal{P}_{t}(\pi)=\mu_{t},\quad t\in[T].$ (7b)
@@
A coupling of $(\mu_{0},\dots,\mu_{T})$ that attains the minimum in problem (7) is said to be an optimal coupling or an optimal transport plan. Problem (7) can be generalised by relaxing the constraint (7b) to only hold for a subset $\Gamma\subseteq[T]$ of the marginals or by introducing unbalanced problems [4, 17, 26] (see also [31, 48, 49, 58] for bi-marginal problems).
@@
In order to solve problems on the form (7) computationally, we assume that each of the sets $\mathcal{X}_{0},\dots,\mathcal{X}_{T}$ is finite, that is, $\cup_{t\in[T]}\mathcal{X}_{t}$ is contained within $n\in\mathbb{N}$ points. Then the marginals can be represented by vectors $m_{t}\in\mathbb{R}^{n}$, $t\in[T]$, and the cost function and the transport plan by tensors $\mathbf{C}\in\mathbb{R}^{n^{T+1}}$ and $\mathbf{Q}\in\mathbb{R}_{+}^{n^{T+1}}$, respectively. The discrete analogue to problem (7) is then
$\underset{\mathbf{Q}\in\mathbb{R}_{+}^{n^{T+1}}}{\min}\quad$ $\langle\mathbf{C},\mathbf{Q}\rangle$ (8a)
subject to $P_{t}(\mathbf{Q})=m_{t},\quad t\in[T].$ (8b)
@@
Problem (8) is a linear programming problem in $n^{T+1}$ variables. Since the number of variables grows exponentially as a function of the number of marginals, the problem quickly becomes intractable for standard methods — even problems with two marginals ($T=1$) are not straightforward to solve when the number of points $n$ is large. One way to alleviate this difficulty is to regularise the problem with an entropy term, as proposed in [9, 19, 57]; we will see that this leads to a dual problem in $n(T+1)$ variables and explicit formulas for the optimal choice thereof.
@@
$\underset{\mathbf{Q}\in\mathbb{R}_{+}^{n^{T+1}}}{\min}\quad$ $\langle\mathbf{C},\mathbf{Q}\rangle+\varepsilon D(\mathbf{Q})$ (9a)
subject to $P_{t}(\mathbf{Q})=m_{t},\quad t\in[T],$ (9b)
@@
$D(\mathbf{Q}):=\sum_{i_{0},\dots,i_{T}\in[n-1]}\mathbf{Q}(i_{0},\dots,i_{T})\log\mathbf{Q}(i_{0},\dots,i_{T})-\mathbf{Q}(i_{0},\dots,i_{T})$
@@
$\mathcal{L}(\mathbf{Q},\lambda)=\langle\mathbf{C},\mathbf{Q}\rangle+\varepsilon D(\mathbf{Q})+\sum_{t\in[T]}\lambda_{t}^{\top}\left(m_{t}-P_{t}(\mathbf{Q})\right),$
@@
where $\lambda=\{\lambda_{t}\}_{t\in[T]}$ are Lagrangian multipliers in $\mathbb{R}^{n}$ corresponding to the constraints (9b). The Lagrangian multipliers are also referred to as dual variables, while the dual functional of problem (9) is defined as $\varphi(\lambda):=\min\{\mathcal{L}(\mathbf{Q},\lambda):\mathbf{Q}\in\mathbb{R}_{+}^{n^{T+1}}\}$.
@@
Since the regularisation term $D(\mathbf{Q})$ ensures that the Lagrangian function is strictly convex in $\mathbf{Q}$, the minimum is attained by the unique $\mathbf{Q}>0$ for which the gradient of the Lagrangian function is zero; that is, for the transport plan $\mathbf{Q}$ satisfying
$0=\partial_{\mathbf{Q}(i_{0},\ldots,i_{T})}\mathcal{L}(\mathbf{Q},\lambda)=\mathbf{C}(i_{0},\ldots,i_{T})+\varepsilon\log\mathbf{Q}(i_{0},\ldots,i_{T})-\sum_{t\in[T]}\lambda_{t}(i_{t})$
@@
for $i_{0},\dots,i_{T}\in\{1,\dots,n\}$. Solving for $\mathbf{Q}$ yields that the stationary point is given by $\mathbf{Q}^{\lambda}=\mathbf{K}\odot\mathbf{U}^{\lambda}$, where
$\mathbf{K}(i_{0},\dots,i_{T}):=e^{-\mathbf{C}(i_{0},\dots,i_{T})/\varepsilon},\quad\mathbf{U}^{\lambda}(i_{0},\dots,i_{T}):=\prod_{t\in[T]}e^{\lambda_{t}(i_{t})/\varepsilon},\quad i_{0},\dots,i_{T}=1,\dots,n.$